### v2

@@ B.4.4 Illustrative Example
An illustrative example is shown in Tab. 15. All samples for T4 are publicly available.
@@
The model prompt is shown in Tab. 16.
@@ B.4.6 Checklist-mapped Reference
Instead, to evaluate the quality of the generated output according to the rubric , we extracted comparable information from the reference tutor data. Section B.4.3 already describes how to identify elements that describe a problem and corresponding solutions. Additionally, it provides a method for determining whether the student is struggling with the current sub-problem—specifically by checking for repeated errors across attempts. This Repeated mistake attribute is used to assign appropriate values to the Error Identification type and Error Rectification type. These types are set to Indirect when the student is attempting the sub-problem for the first time. If the Repeated mistake flag is active, both types are instead set to Direct. Notably, these properties are only assigned when the student’s response contains an error. For the Sub-question Transition property, we directly use the value provided in the original dataset, which is only set when the student gives a correct answer but the overall problem remains unsolved. An example of the checklist-mapped human reference from Tab. 15 is shown in Tab. 17.
@@ B.5.2 Task Significance
Our motivation to assess LLMs’ ability to evaluate student responses beyond student-facing rubric stems from the practical insight that instructors often do not rely solely on a rubric Instead, they apply holistic judgments that incorporate broader, often implicit, expectations [60, 61]. This is particularly evident in writing assignments, where public rubrics tend to be too general to support nuanced evaluation. Instructors frequently augment them with private, instance-specific criteria to enable more accurate and context-sensitive scoring [62]. Building on these observations, we seek to determine whether LLMs can similarly move beyond surface-level guidelines to produce evaluations that align with instructor judgment.
@@ B.5.3 Data Acquisition
For this task, we use data from the Economics 101 course at the University of Michigan, Ann Arbor [63]. The assignment presents a scenario in which “an increase in the minimum wage in San Francisco could lead to increased adoption of automation.” To counter this, two policy options are proposed: (a) a tax on automation and (b) a ban on automation. Students are asked to write a persuasive letter outlining the economic implications of the wage increase and to argue against one of the proposed policies using concepts and tools covered in the course. The student-facing rubric is shown in Tab. 18 - it is clearly evident that the student-facing rubric is high-level and cannot be used as a criteria for robustly assigning the scores.
@@ B.5.4 Illustrative Example
Due to the proprietary nature of the data, we cannot provide a specific example. However, an example of instructor feedback is provided in Tab. 19 for a student written essay and the model prompt is shown in Tab. 20:
@@ B.5.6 Checklist-mapped Reference
A prompt for creating the checklist-mapped reference is not necessary in this task, as all model responses are binary (“Yes” or “No”) and easily verifiable. An example of the checklist-mapped reference of the instructor feedback in Tab. 19 is presented in Tab. 21.
@@ B.6.1 Task Definition
Clinical note generation is the task of producing well-structured, accurate, and comprehensive clinical notes based on patient-doctor dialogue during a clinical encounter [64]. In this task, we focus on generating SOAP notes, as they are the most common way to document medical interactions [65]. SOAP notes consist of four key sections: Subjective (patient-reported symptoms), Objective (clinician-observed data), Assessment (diagnoses or impressions), and Plan (treatment or follow-up steps). The objective is to generate accurate, structured, and useful SOAP notes from unstructured patient and doctor conversations. Automating this process would greatly reduce the clinician workload and enhance the quality of medical documentation. In using LLMs to perform this task, we input a the transcript of a conversation between a healthcare professional and patient. The output is a SOAP note that includes the key information from the conversation. An example note is shown in Tab. 22.
@@ B.6.2 Task Significance
Clinical documentation plays an important role in effective, high-quality care by supporting accurate diagnoses and standardizing communication among healthcare providers. However, the generation of structured clinical notes is time-consuming and prone to human error. Studies show that physicians can spend over an hour documenting a single clinical visit and nearly two hours on electronic health record (EHR) tasks for every hour of direct patient care [66]. Given the demanding schedules of healthcare professionals, this burden increases the risk of misremembered or inaccurately entered information. Automating this process with large language models (LLMs) can improve consistency and streamline workflows, ultimately reducing patient wait times and enhancing the quality of care.
@@ B.6.3 Data Acquisition
We use the ACI-Bench dataset, a benchmark for clinical note generation with 207 annotated clinical encounters. Each sample includes a transcript of a patient-doctor interaction and its corresponding SOAP structured note written by the doctor [67]. Transcripts were created using three methods: (1) a virtual assistant, in which the doctor had to use explicit terms (e.g. “Hey Dragon show me the diabetes lab”) during the visit; (2) a virtual scribe, automated or otherwise, which assisted in note creation without distracting in-person doctor-patient interactions; and (3) an ambient clinical intelligence (ACI), which created the transcript without interrupting natural conversation flow between the patient and physician.
From these 207 samples, we selected a final subset of 100 on difficulty and diversity consistent with the rationale in §3.2. Difficulty was measured using the length of the human reference. Longer SOAP notes will contain more information about the patient, treatment, and analysis, making diagnosis a more difficult task. Additionally, ACI-Bench was originally collected in a diverse manner. Not only were transcripts created in a variety of methods, the content of the SOAP note itself was also diverse. For example, some notes contained Chief Complaints, but others did not; other notes had longer History of Present Illness sections, while others included symptom context in the Assessment or Plan sections.
@@ B.6.4 Illustrative Example
An illustrative example is shown in Tab. 22. All samples for T6 are publicly available.
@@
The model prompt shown in Tab. 23 was taken directly from the ACI-Bench benchmark [67]. Although the prompt does not explicitly specify the four SOAP sections, the paper explains that it was strongly adapted from the SOAP note format: the History of Present Illness aligns with the Subjective section, the Exam and Results correspond to the Objective section, and the Assessment and Plan are combined into the final section. This structure also justifies our use of checklist items that incorporate SOAP elements in Section B.6.5.
@@ B.6.5 Evaluation Rubric
In T6, we did not consult domain experts as in previous tasks, but created our rubric independently through extensive research. Our rubric items and definitions were verified across multiple reputable online sources [64, 65] and are still expert-level. The rubric comprises 29 checklist items that captures the key information present in a SOAP note. The detailed checklist items are listed as follows:
@@ B.6.6 Checklist-mapped Reference
The checklist-mapped references for both the human reference and model output were constructed with a similar approach to that used in T1 (see Section B.1.6). However, in T6, we grouped 29 items into five groups based on (1) the expected length of the model’s response and (2) the SOAP category the item belonged to (e.g., Subjective, Objective, etc.). This is to help the model focus on distinct clinical sections and break down complex notes into meaningful, manageable parts. The prompt for extracting an individual checklist item is presented in Tab. 24. A checklist-mapped reference of the sample in Tab. 22 is shown in Tab. 25.
To assess the quality of the checklist-mapped reference, we selected an additional 30 difficult and diverse samples, different from the main dataset, and conducted human verification to examine the faithfulness of the model response. Details of this experiment can be found in LABEL:ap:validation-of-checklist.
@@ B.7.1 Task Definition
Molecule description generation involves creating accurate and structured natural language descriptions of molecular structures based on their SMILES (Simplified Molecular Input Line Entry System) representations [68]. SMILES encodes molecules as linear strings that represent atoms, bonds, rings, and branching patterns, and serves as a textual representation of molecular graphs. The objective is to translate these symbolic sequences into natural language descriptions that capture key structural and chemical features of the molecule. When using LLMs to complete this task, the input consists of a SMILES string (e.g., “CC(=O)OC1=CC=CC=C1C(=O)O”). The output is a molecule description a professional might read. An example is shown in Tab. 26.
@@ B.7.2 Task Significance
Chemical databases contain tens of millions of molecules, each represented by complex SMILES strings that are not easily interpretable by humans. Experts often find it time-consuming and challenging to infer the functional class or properties of a molecule solely from its SMILES representation, especially for more complex structures. Automating the translation of SMILES into structured, ontological, and natural language descriptions can significantly enhance the accessibility and usability of chemical data. This advancement supports key fields such as drug discovery, materials science, and chemical education by enabling quicker understanding and communication of molecular information [69].
For example, models such as MolT5 have been specifically developed to complete this task. MolT5 leverages transformer-based architectures to improve the quality and informativeness of molecular descriptions, facilitating a deeper understanding of molecular structures and their properties [70]. Such models not only accelerate the discovery phase but also improve collaboration across multidisciplinary research teams.
@@ B.7.3 Data Acquisition and Preprocessing
We use the ChEBI-20 dataset originally collected for the Text2Mol task [69]. This dataset was created by collecting compound annotations from the ChEBI database [^20] https://www.ebi.ac.uk/chebi/, which were scraped from PubChem [^21] https://pubchem.ncbi.nlm.nih.gov/. Descriptions shorter than 20 words were excluded to ensure sufficient detail. The resulting dataset contains 33,010 pairs of SMILES strings and their corresponding textual descriptions. Both ChEBI and PubChem are specialized, domain-specific chemical resources commonly utilized by chemists, aligning well with the expert-level knowledge demanded by this task.
@@ B.7.4 Illustrative Example
An illustrative example is shown in Tab. 26. All samples for T7 are publicly available.
@@
The model prompt is presented in Tab. 27. This prompt was carefully designed in collaboration with domain experts to ensure both generality and relevance across a broad range of chemical compounds. To align with the ChEBI dataset, which stands for “Chemical Entities of Biological Interest,” we incorporated the phrase “Chemicals of Biological significance” into the prompt. However, we deliberately avoided including the acronym “ChEBI” itself to prevent the model from relying on any dataset-specific shortcuts.
@@ B.7.5 Evaluation Rubric
The final rubric comprises of six checklist items that capture key information typically found in professional molecule descriptions. The last sentence of each definition (starting with “Usually, but not always,”) is included to support the construction of the checklist-mapped reference described in Section B.7.6. The detailed checklist items are listed as follows:
@@ B.7.6 Checklist-mapped Reference
In line with the method described in T1, we generated checklist-mapped references for the human reference and model outputs with our evaluation rubric. However, with only six checklist items, we were able to create one group with one model extraction prompt. The prompt for extracting an individual checklist item is presented in Tab. 28 and the model output extraction prompt is available in our public GitHub repository. The checklist-mapped reference for the sample in Tab. 26 is shown in Tab. 29.
@@ B.8.1 Task Definition
Protein description generation, also known as protein captioning, refers to creating an accurate, informative, and useful description of a protein given its amino acid sequence. These sequences are composed of letters representing individual amino acids and are the standard linear representations of protein primary structure. The objective is to generate a natural language description that captures essential biological characteristics of the protein, such as its function, cellular location, family or domain classification, and any relevant structural or catalytic properties. An example of a sequence-description pair is shown in Tab. 30. In using LLMs to complete this task, the input is an amino acid sequence (e.g., MKWVTFISLLFLFSSAYSRGVFRRDTH…) and the output is a protein description [71].
@@ B.8.2 Task Significance
Protein sequences are long, complex, and typically require expert analysis and database lookups (e.g., UniProt [^22] https://www.uniprot.org, PDB [^23] https://www.rcsb.org/) to determine their biological functions. Automating the generation of descriptions helps make protein information more accessible and supports tasks like genome annotation, database curation, and bioinformatics research. Recent tools such as AnnoPRO [72] apply deep learning to predict protein function from sequences, reducing the need for manual curation and accelerating biological discovery [72].
@@ B.8.3 Data Acquisition and Preprocessing
We collected protein sequences and their corresponding descriptions from the SciKnowEval dataset, which sources its protein and caption entries from the UniProtKB database [^24] https://www.uniprot.org/help/uniprotkb —a comprehensive, curated resource for protein sequence and functional information [73]. Reference descriptions are written by experts, curated by UniProt, and serve as the human reference for this task. From this dataset, we filtered a subset of 100 high-quality samples for our task. These samples were selected on length, all of which are over 900 characters.
@@ B.8.4 Illustrative Example
An illustrative example is shown in Tab. 30 All samples for T8 are publicly available.
@@
Considering contextual grounding, we adopt a role-based prompting strategy. The model is instructed to assume the role of a protein researcher tasked with writing descriptions for protein sequences and encouraged to use domain-specific terminology with the model prompt shown in Tab. 31.
@@ B.8.5 Evaluation Rubric
We collaborated with two graduate students from the University of Michigan studying biology to create the evaluation rubric. We first presented an initial draft of the rubric, created using online sources, and fine-tuned it with the experts’ assistance across three separate meetings until both experts agreed with the rubric. The experts’ also provided us with sources that explained their reasoning for including or removing certain items in the rubric. The rubric comprises five checklist items that captures the key information present in a protein description. The detailed checklist items are listed as follows:
@@
    Domains/Motifs: Functional regions (e.g., kinase domain, zinc finger) [74, 75].
@@
      Functional Role: How a molecule contributes to biological systems, including the processes it is involved in (e.g., metabolism, cellular functions) and the specific role it plays (e.g., enzymatic, structural, or signaling) [76, 77].
@@
    Interactions: The physical and functional associations within a cell or organism, including but not limited to protein partners, ligands/substrates, and cofactors [76, 77].

### v3

@@ B.4.4 Illustrative Example
An illustrative example is shown in Tab.˜15. All samples for T4 are publicly available.
@@
The model prompt is shown in Tab.˜16.
@@ B.4.6 Checklist-mapped Reference
Instead, to evaluate the quality of the generated output according to the rubric , we extracted comparable information from the reference tutor data. Section˜B.4.3 already describes how to identify elements that describe a problem and corresponding solutions. Additionally, it provides a method for determining whether the student is struggling with the current sub-problem—specifically by checking for repeated errors across attempts. This Repeated mistake attribute is used to assign appropriate values to the Error Identification type and Error Rectification type. These types are set to Indirect when the student is attempting the sub-problem for the first time. If the Repeated mistake flag is active, both types are instead set to Direct. Notably, these properties are only assigned when the student’s response contains an error. For the Sub-question Transition property, we directly use the value provided in the original dataset, which is only set when the student gives a correct answer but the overall problem remains unsolved. An example of the checklist-mapped human reference from Tab.˜15 is shown in Tab.˜17.
@@ B.5.2 Task Significance
Our motivation to assess LLMs’ ability to evaluate student responses beyond student-facing rubric stems from the practical insight that instructors often do not rely solely on a rubric Instead, they apply holistic judgments that incorporate broader, often implicit, expectations [Bloxham et al., 2011, Jeong, 2015]. This is particularly evident in writing assignments, where public rubrics tend to be too general to support nuanced evaluation. Instructors frequently augment them with private, instance-specific criteria to enable more accurate and context-sensitive scoring [Broad, 2003]. Building on these observations, we seek to determine whether LLMs can similarly move beyond surface-level guidelines to produce evaluations that align with instructor judgment.
@@ B.5.3 Data Acquisition
For this task, we use data from the Economics 101 course at an R1 university in the US [Nair et al., 2024]. The assignment presents a scenario in which “an increase in the minimum wage in San Francisco could lead to increased adoption of automation.” To counter this, two policy options are proposed: (a) a tax on automation and (b) a ban on automation. Students are asked to write a persuasive letter outlining the economic implications of the wage increase and to argue against one of the proposed policies using concepts and tools covered in the course. The student-facing rubric is shown in Tab.˜18 - it is clearly evident that the student-facing rubric is high-level and cannot be used as a criteria for robustly assigning the scores.
@@ B.5.4 Illustrative Example
Due to the proprietary nature of the data, we cannot provide a specific example. However, an example of instructor feedback is provided in Tab.˜19 for a student written essay and the model prompt is shown in Tab.˜20:
@@ B.5.6 Checklist-mapped Reference
A prompt for creating the checklist-mapped reference is not necessary in this task, as all model responses are binary (“Yes” or “No”) and easily verifiable. An example of the checklist-mapped reference of the instructor feedback in Tab.˜19 is presented in Tab.˜21.
@@ B.6.1 Task Definition
Clinical note generation is the task of producing well-structured, accurate, and comprehensive clinical notes based on patient-doctor dialogue during a clinical encounter [Contributors, 2023]. In this task, we focus on generating SOAP notes, as they are the most common way to document medical interactions Podder et al. [2023]. SOAP notes consist of four key sections: Subjective (patient-reported symptoms), Objective (clinician-observed data), Assessment (diagnoses or impressions), and Plan (treatment or follow-up steps). The objective is to generate accurate, structured, and useful SOAP notes from unstructured patient and doctor conversations. Automating this process would greatly reduce the clinician workload and enhance the quality of medical documentation. In using LLMs to perform this task, we input a the transcript of a conversation between a healthcare professional and patient. The output is a SOAP note that includes the key information from the conversation. An example note is shown in Tab.˜22.
@@ B.6.2 Task Significance
Clinical documentation plays an important role in effective, high-quality care by supporting accurate diagnoses and standardizing communication among healthcare providers. However, the generation of structured clinical notes is time-consuming and prone to human error. Studies show that physicians can spend over an hour documenting a single clinical visit and nearly two hours on electronic health record (EHR) tasks for every hour of direct patient care Arndt et al. [2017]. Given the demanding schedules of healthcare professionals, this burden increases the risk of misremembered or inaccurately entered information. Automating this process with large language models (LLMs) can improve consistency and streamline workflows, ultimately reducing patient wait times and enhancing the quality of care.
@@ B.6.3 Data Acquisition
We use the ACI-Bench dataset, a benchmark for clinical note generation with 207 annotated clinical encounters. Each sample includes a transcript of a patient-doctor interaction and its corresponding SOAP structured note written by the doctor [Yim et al., 2023]. Transcripts were created using three methods: (1) a virtual assistant, in which the doctor had to use explicit terms (e.g. “Hey Dragon show me the diabetes lab”) during the visit; (2) a virtual scribe, automated or otherwise, which assisted in note creation without distracting in-person doctor-patient interactions; and (3) an ambient clinical intelligence (ACI), which created the transcript without interrupting natural conversation flow between the patient and physician.
From these 207 samples, we selected a final subset of 100 on difficulty and diversity consistent with the rationale in §3. Difficulty was measured using the length of the human reference. Longer SOAP notes will contain more information about the patient, treatment, and analysis, making diagnosis a more difficult task. Additionally, ACI-Bench was originally collected in a diverse manner. Not only were transcripts created in a variety of methods, the content of the SOAP note itself was also diverse. For example, some notes contained Chief Complaints, but others did not; other notes had longer History of Present Illness sections, while others included symptom context in the Assessment or Plan sections.
@@ B.6.4 Illustrative Example
An illustrative example is shown in Tab.˜22. All samples for T6 are publicly available.
@@
The model prompt shown in Tab.˜23 was taken directly from the ACI-Bench benchmark Yim et al. [2023]. Although the prompt does not explicitly specify the four SOAP sections, the paper explains that it was strongly adapted from the SOAP note format: the History of Present Illness aligns with the Subjective section, the Exam and Results correspond to the Objective section, and the Assessment and Plan are combined into the final section. This structure also justifies our use of checklist items that incorporate SOAP elements in Section˜B.6.5.
@@ B.6.5 Evaluation Rubric
In T6, we did not consult domain experts as in previous tasks, but created our rubric independently through extensive research. Our rubric items and definitions were verified across multiple reputable online sources [Contributors, 2023, Podder et al., 2023] and are still expert-level. The rubric comprises 29 checklist items that captures the key information present in a SOAP note. The detailed checklist items are listed as follows:
@@ B.6.6 Checklist-mapped Reference
The checklist-mapped references for both the human reference and model output were constructed with a similar approach to that used in T1 (see Section˜B.1.6). However, in T6, we grouped 29 items into five groups based on (1) the expected length of the model’s response and (2) the SOAP category the item belonged to (e.g., Subjective, Objective, etc.). This is to help the model focus on distinct clinical sections and break down complex notes into meaningful, manageable parts. The prompt for extracting an individual checklist item is presented in Tab.˜24. A checklist-mapped reference of the sample in Tab.˜22 is shown in Tab.˜25.
To assess the quality of the checklist-mapped reference, we selected an additional 30 difficult and diverse samples, different from the main dataset, and conducted human verification to examine the faithfulness of the model response. Details of this experiment can be found in Section˜E.1.
@@ B.7.1 Task Definition
Molecule description generation involves creating accurate and structured natural language descriptions of molecular structures based on their SMILES (Simplified Molecular Input Line Entry System) representations [Weininger, 1988]. SMILES encodes molecules as linear strings that represent atoms, bonds, rings, and branching patterns, and serves as a textual representation of molecular graphs. The objective is to translate these symbolic sequences into natural language descriptions that capture key structural and chemical features of the molecule. When using LLMs to complete this task, the input consists of a SMILES string (e.g., “CC(=O)OC1=CC=CC=C1C(=O)O”). The output is a molecule description a professional might read. An example is shown in Tab.˜26.
@@ B.7.2 Task Significance
Chemical databases contain tens of millions of molecules, each represented by complex SMILES strings that are not easily interpretable by humans. Experts often find it time-consuming and challenging to infer the functional class or properties of a molecule solely from its SMILES representation, especially for more complex structures. Automating the translation of SMILES into structured, ontological, and natural language descriptions can significantly enhance the accessibility and usability of chemical data. This advancement supports key fields such as drug discovery, materials science, and chemical education by enabling quicker understanding and communication of molecular information [Edwards et al., 2021].
For example, models such as MolT5 have been specifically developed to complete this task. MolT5 leverages transformer-based architectures to improve the quality and informativeness of molecular descriptions, facilitating a deeper understanding of molecular structures and their properties [Edwards et al., 2022]. Such models not only accelerate the discovery phase but also improve collaboration across multidisciplinary research teams.
@@ B.7.3 Data Acquisition and Preprocessing
We use the ChEBI-20 dataset originally collected for the Text2Mol task [Edwards et al., 2021]. This dataset was created by collecting compound annotations from the ChEBI database [^20] https://www.ebi.ac.uk/chebi/, which were scraped from PubChem [^21] https://pubchem.ncbi.nlm.nih.gov/. Descriptions shorter than 20 words were excluded to ensure sufficient detail. The resulting dataset contains 33,010 pairs of SMILES strings and their corresponding textual descriptions. Both ChEBI and PubChem are specialized, domain-specific chemical resources commonly utilized by chemists, aligning well with the expert-level knowledge demanded by this task.
@@ B.7.4 Illustrative Example
An illustrative example is shown in Tab.˜26. All samples for T7 are publicly available.
@@
The model prompt is presented in Tab.˜27. This prompt was carefully designed in collaboration with domain experts to ensure both generality and relevance across a broad range of chemical compounds. To align with the ChEBI dataset, which stands for “Chemical Entities of Biological Interest,” we incorporated the phrase “Chemicals of Biological significance” into the prompt. However, we deliberately avoided including the acronym “ChEBI” itself to prevent the model from relying on any dataset-specific shortcuts.
@@ B.7.5 Evaluation Rubric
The final rubric comprises of six checklist items that capture key information typically found in professional molecule descriptions. The last sentence of each definition (starting with “Usually, but not always,”) is included to support the construction of the checklist-mapped reference described in Section˜B.7.6. The detailed checklist items are listed as follows:
@@ B.7.6 Checklist-mapped Reference
In line with the method described in T1, we generated checklist-mapped references for the human reference and model outputs with our evaluation rubric. However, with only six checklist items, we were able to create one group with one model extraction prompt. The prompt for extracting an individual checklist item is presented in Tab.˜28 and the model output extraction prompt will be available in our public GitHub repository. The checklist-mapped reference for the sample in Tab.˜26 is shown in Tab.˜29.
@@ B.8.1 Task Definition
Protein description generation, also known as protein captioning, refers to creating an accurate, informative, and useful description of a protein given its amino acid sequence. These sequences are composed of letters representing individual amino acids and are the standard linear representations of protein primary structure. The objective is to generate a natural language description that captures essential biological characteristics of the protein, such as its function, cellular location, family or domain classification, and any relevant structural or catalytic properties. An example of a sequence-description pair is shown in Tab.˜30. In using LLMs to complete this task, the input is an amino acid sequence (e.g., MKWVTFISLLFLFSSAYSRGVFRRDTH…) and the output is a protein description [Rives et al., 2021].
@@ B.8.2 Task Significance
Protein sequences are long, complex, and typically require expert analysis and database lookups (e.g., UniProt [^22] https://www.uniprot.org, PDB [^23] https://www.rcsb.org/) to determine their biological functions. Automating the generation of descriptions helps make protein information more accessible and supports tasks like genome annotation, database curation, and bioinformatics research. Recent tools such as AnnoPRO Li et al. [2023b] apply deep learning to predict protein function from sequences, reducing the need for manual curation and accelerating biological discovery [Li et al., 2023b].
@@ B.8.3 Data Acquisition and Preprocessing
We collected protein sequences and their corresponding descriptions from the SciKnowEval dataset, which sources its protein and caption entries from the UniProtKB database [^24] https://www.uniprot.org/help/uniprotkb —a comprehensive, curated resource for protein sequence and functional information [Feng et al., 2024]. Reference descriptions are written by experts, curated by UniProt, and serve as the human reference for this task. From this dataset, we filtered a subset of 100 high-quality samples for our task. These samples were selected on length, all of which are over 900 characters.
@@ B.8.4 Illustrative Example
An illustrative example is shown in Tab.˜30 All samples for T8 are publicly available.
@@
Considering contextual grounding, we adopt a role-based prompting strategy. The model is instructed to assume the role of a protein researcher tasked with writing descriptions for protein sequences and encouraged to use domain-specific terminology with the model prompt shown in Tab.˜31.
@@ B.8.5 Evaluation Rubric
We collaborated with two graduate students from an R1 university at the US studying biology to create the evaluation rubric. We first presented an initial draft of the rubric, created using online sources, and fine-tuned it with the experts’ assistance across three separate meetings until both experts agreed with the rubric. The experts’ also provided us with sources that explained their reasoning for including or removing certain items in the rubric. The rubric comprises five checklist items that captures the key information present in a protein description. The detailed checklist items are listed as follows:
@@
    Domains/Motifs: Functional regions (e.g., kinase domain, zinc finger) [Murzin et al., 1995, Berman et al., 2000].
@@
      Functional Role: How a molecule contributes to biological systems, including the processes it is involved in (e.g., metabolism, cellular functions) and the specific role it plays (e.g., enzymatic, structural, or signaling) [Szklarczyk et al., 2023, Oughtred et al., 2021].
@@
    Interactions: The physical and functional associations within a cell or organism, including but not limited to protein partners, ligands/substrates, and cofactors [Szklarczyk et al., 2023, Oughtred et al., 2021].